You are caring for a client that is on a continuous heparin infusion. The order of the infusion is the following heparin IV 18units/kg/hour; available heparin is 25,000units/500ml; the client weighs 240pounds. You are giving your nurse report at the end of you shift when the nurse taking over care for the client says, "is the heparin supposed to be infusing at 49ml/hr?" You double check the rate and the pump is set to 49ml/hr. What are your top 3 priorities (include your rationale)?
Answer #1

Ordered dose of heparin is 18units /kg/hour

Available heparin is 25000units / 500ml

Weight of the client is 240 pounds = 240 × 0.45= 108kg

Then ordered dose will be 108× 18 = 1944 units /hour

Flow rate = ordered dose / available dose × volume

= 1944/ 25000 × 500

= 38.8ml /hr = 39 ml/hr should be the flow rate

Top 3 priorities are :-

- inform the physician about increased dose of heparin given

- change the dose as per advice

- take necessary action as per doctors advice to avoid complications related to increased dosage of heparin